### v1

# DreamAudio: Customized Text-to-Audio Generation with Diffusion Models

Abstract
With the development of large-scale diffusion-based and language-modeling-based generative models, impressive progress has been achieved in text-to-audio generation. Despite producing high-quality outputs, existing text-to-audio models mainly aim to generate semantically aligned sound and fall short on precisely controlling fine-grained acoustic characteristics of specific sounds. As a result, users that need specific sound content may find it challenging to generate the desired audio clips. In this paper, we present DreamAudio for customized text-to-audio generation (CTTA). Specifically, we introduce a new framework that is designed to enable the model to identify auditory information from user-provided reference concepts for audio generation. Given a few reference audio samples containing personalized audio events, our system can generate new audio samples that include these specific events. In addition, two types of datasets are developed for training and testing the customized systems. The experiments show that the proposed model, DreamAudio, generates audio samples that are highly consistent with the customized audio features and aligned well with the input text prompts. Furthermore, DreamAudio offers comparable performance in general text-to-audio tasks. We also provide a human-involved dataset containing audio events from real-world CTTA cases as the benchmark for customized generation tasks. Demos are available at https://yyua8222.github.io/DreamAudio_demopage/.
Index Terms:
audio generation, diffusion model, retrieval argumentation, customized generation, AIGC
I Introduction
Audio generation, as a crucial technology for enabling artificial intelligence generated content (AIGC) [1], has gained significant interest from the research community. In recent years, conditional audio generation has become a popular paradigm, where music, speech, or general sound effects are generated based on a variety of conditions, such as text [2], images [3, 4], and videos [5, 6]. This opened new opportunities for a range of potential applications, including audio generation for movies [7], games [8], and audiobooks [9].
Facilitated by advances in diffusion-based generative models [10, 11] and large-scale audio-language datasets [12, 13, 14, 15], several models have been developed for audio generation with text prompts as conditions, such as AudioLDM [2], AudioGen [16], DiffSound [17], TANGO [18], Make-an-Audio2 [19], Re-AudioLDM [20], and AudioLDM2 [21]. Building on the “semantic prior” learned from large collections of datasets [22] to associate textual concepts with audio features, these models have shown strong capabilities in generating audio samples of high quality, fidelity and diversity. For example, when provided with the text prompt dog barking, the model leverages this particular “semantic prior” to generate various audio clips depicting dog barking across varying species, emotions, and durations.
Despite significant progress in diffusion-based generative systems [23, 24, 25, 26, 27], current text-to-audio generation systems often lack the flexibility to customize content based on personalized intentions, particularly when dealing with rare or unseen sounds. This can cause problems in real-world applications for multimedia content production which often requires the generation of audio samples to be tailored for specific features. For example, it can be challenging for current TTA models to generate sounds that are consistent with the text prompt “a monster is fighting with a Minion”. This is because such audio events, like “monster fight” and “Minion talk” are unique and rare, or with a specific timbre, which can hardly be found from any current training data. Hence, users often need to go through several rounds of trial and error to produce the desired output and may encounter difficulties in achieving the optimal result [28].
Figure 1: A demo of DreamAudio for audio generation with customized content of “monster fighting” and “Minion talking”. The system takes both the text prompt and user-provided audio-caption pairs as the reference concepts, and generate audio content consistent with the description ”Monster is fighting with a Minion”.
To address this challenge, several works have explored the ideas for generating rare or unseen audio events, especially for few-shot or zero-shot scenarios. One of the pioneering works is the Re-AudioLDM [20], which is a retrieval-based technique to improve the performance for the generation of rare audio events using audio-caption pairs retrieved from external data. Similarly, AudioBox TTA-RAG [29] introduced a retrieval embedding module to improve the performance for few-shot and zero-shot audio generation tasks. Although these previous approaches effectively improve the semantic accuracy in generating infrequent or unseen audio events, they are still unable to provide explicit control over the sound effects generated, as specified by users.
In this paper, we propose a new TTA task, namely, customized text-to-audio generation (CTTA), where the content produced by the generation system can be controlled and customized by users. For instance, the system is capable to generate the dog barking sound from a specific dog with unique timber, or the sound for a specific monster which is not included in the traning dataset. We refer to these featured audio events as user reference concepts. To achieve this, we introduce DreamAudio, a latent diffusion-based system with flow matching for the CTTA task. More specifically, inspired by ControlNet [30], we propose a multi-reference customization (MRC) structure for the generator module to control the content generated by the model. We design a new group of encoder blocks to extract features from the user-provided reference concepts. Then, the system is trained to fuse such customized features for the generation of the audio output. With a new cross-attention module linking the provided audio concepts with their corresponding language descriptions, our proposed method provides a training-free pipeline that is capable of generating audio samples, consistent with the reference concepts. Figure 1 shows an example of CTTA, where the model is enabled to generate the audio sample with customized “monster fighting” and “Minion talking”, in terms of the provided reference concepts.
To facilitate model development and evaluation, we create two datasets with different formats by concatenating and overlapping different audio events, called Customized-Concatenation and Customized-Overlay. In addition, we collect various special audio events for customized cases, such as fiction and comic artificial sounds, and create a small-scale dataset that better aligns with real-world multimedia production scenarios as a benchmark for CTTA. We perform experiments on these two datasets, and the results show that our method achieves state-of-the-art (SOTA) results in terms of audio generation with customized content. Furthermore, DreamAudio presents comparable results to previous SOTA models on general text-to-audio generation tasks using AudioCaps [31], illustrating substantial advancements in customized TTA tasks and competitive performance in general TTA tasks.
Our contributions can be summarized as follows.
  • We propose a novel audio generation model, DreamAudio, that is capable of performing content-customized audio generation with text prompts as condition.
  • We propose a multi-reference customization structure, which allows the features from the reference audio to be fused for text-to-audio generation.
  • We develop two new datasets for training and evaluating the proposed method and establish a new benchmark for content-customized text-to-audio generation from real-world audio events.
  • Our experiments show that DreamAudio significantly enhances the ability of TTA models for customized audio generation and achieves state-of-the-art CTTA performance.
II Related Work
II-A Diffusion Models
Diffusion-based models [11, 32] have demonstrated improved performance in generative tasks for image [33, 24, 34, 35], audio [36, 37], and video [38, 39]. In the realm of audio synthesis, researchers initially followed the design used in image generation, and adapted it for audio generation, based on mel-spectrogram [40, 41] and waveform [42, 43, 44]. Due to the involvement of high-dimensional data [45, 46], the generation systems based on waveform and spectrogram representations can be difficult to train and slow to infer. To address this limitation, recent diffusion-based models work in a latent space through encoding pipelines [26]. Specifically, current SOTA systems follow an encoder-decoder framework [47], trained to generate the target features in a latent space, which are then decoded and reconstructed into waveforms through vocoders [48, 49, 50].
II-B Conditional Audio Generation
Controllable generation has emerged as an important area in audio synthesis, enabling users to guide generative models with specific constraints or conditions to produce desired outcomes [30]. Text-driven audio generation has gained significant attention. AudioGen [16] employs a conditional language modeling pipeline that generates audio waveforms directly. AudioLDM [2] employs the Contrastive Language-Audio Pretraining (CLAP) [51] model to derive the embeddings of audio and text, and the use them as conditions to guide the training and inference of the latent diffusion model or audio generation. Specifically, Audio is trained to generate feature representations of target audio within the latent space, followed by a variational autoencoder (VAE) decoder to reconstruct the spectrogram from the latent representation, finally using a vocoder to reconstruct the waveforms from the spectrogram. Make-an-Audio develops a pseudo-prompt enhancement strategy to generate extra audio-caption pairs with large-scale compositions to alleviate the data scarcity problem. Tango [52] uses a structure similar to Audio, but replaces the CLAP model with a FLAN-T5 [53].
Due to the diversity of the training dataset [12, 31, 13], these previous methods allow the model to generate highly diverse samples for the same prompt. However, these models are not designed for CTTA tasks and often struggle to generate content from user preferences [20], which is the problem to be addressed in this paper.
Figure 2: The inference pipeline of the DreamAudio. The input prompt and reference concept are encoded in two paralleled paths through the Flan-T5 Encoder and the reference audio feature is encoded by the VAE Encoder. Along with the noised data $\boldsymbol{z}_{\lambda}$, four inputs are forwarded to the generator with MRC structure to generate the denoised data, followed by the VAE decoder and vocoder to reconstruct the final output.
II-C Customized Content Generation
Current methods for customized content generation are developed dominantly for images, such as [54, 55, 56, 57, 58, 59, 60]. They could be categorized into training-based customization [22, 61, 62, 63], and tailoring-based customization [64, 65, 66, 67, 68, 56, 69]. In training-based methods, the generation models are fine-tuned with the embeddings of specific subjects, in order to bind a unique identifier and control the generated contents. In tailoring-based methods, customization is achieved by integrating the target feature with the generated feature during the inference stages to minimize the fine-tuning processes.
The CTTA task remains to be explored. Most similar to the CTTA task are retrieval-based TTA tasks, developed recently to improve the performance of TTA models in the few-shot and zero-shot cases, such as Re-AudioLDM [20], which incorporates external audio features into the generation process using retrieval-augmented approaches [70, 71]. This system demonstrates more stable performance on low-occurrence audio events in TTA generation. Building on this concept, AudioBox TTA-RAG [29] introduces a retrieval-information embedding module to enhance the capability in zero-shot generation. Although these systems can improve the TTA generation performance in the few-shot and zero-shot cases, the quality of the audio samples is limited by the database for audio retrieval. Furthermore, these models lack the capability to generate specific content for customized tasks and the research for CTTA tasks remains to be explored.
In this paper, we design a novel CTTA system, which allows the generated audio content to be tailored and controlled in terms of user preference derived from the reference concepts, in addition to the text prompts. The ability to achieve customization not only enhances the creative process, but also improves controllability in TTA systems, thereby addressing a crucial practical challenge for deploying TTA systems in real-world applications.
III Proposed Method
III-A Overview of the Proposed Model
Given only a few audio-language reference pairs $\{(\mathbf{a}_{1},\mathbf{t}_{1}),(\mathbf{a}_{2},\mathbf{t}_{2}),\dots,(\mathbf{a}_{K},\mathbf{t}_{K})\}$, where each reference audio clip $\mathbf{a}_{k}$ is associated with a reference caption $\mathbf{t}_{k}$, k=1,,Kk=1,...,K, the CTTA task aims to generate the target audio $\mathbf{a}$ conditioned on the input textual prompt $\boldsymbol{c}$ and the reference concepts, derived from pairs of reference audios $\{\mathbf{a}_{1},...,\mathbf{a}_{K}\}$ with their corresponding captions $\{\mathbf{t}_{1},...,\mathbf{t}_{K}\}$.
To address this task, we propose the DreamAudio system, as illustrated in Figure 2. The system is composed of several modules, including the text encoder (e.g. Flan-T5) to obtain the embeddings of the input text prompt and reference captions, and the audio encoder (e.g. VAE encoder) to obtain the embeddings of the reference audios, which are followed by the feature generator built on rectified flow matching (RFM). To customize the audio features with reference concepts, multi-reference customization (MRC) structure is designed for the feature generator. The customized audio embeddings are then converted to spectrograms using a VAE decoder, which is then turned into a waveform using a vocoder (e.g. BigVGAN [50]).
We first discuss the calculation of text and audio embeddings in Section III-B. Then, we present the implementation of our framework for diffusion-based feature generatorin Section III-C, starting with the preliminaries of module training strategy, followed by the details of the proposed MRC architecture. Finally, Section III-D introduces the process for reconstructing the target audio feature.
III-B Text and Audio Embeddings
III-B1 Text Embedding
[…]
$\boldsymbol{C}=\textit{f}_{\text{T5}}(\boldsymbol{c})$ (1)
where $\textit{f}_{\text{T5}}(\cdot)$ is a text encoder, chosen to be the FLAN-T5 model [53] in this work.
For each caption of the reference audio concepts, $\boldsymbol{t}_{k},k=1,...,K$, we first use the FLAN-T5 model to compute its embedding as $\boldsymbol{e}_{k}=\textit{f}_{\text{T5}}(\boldsymbol{t}_{k})$. We then concatenate the embeddings of all the reference captions as the second input for the module, as follows,
$\boldsymbol{E}=[\boldsymbol{e}_{1},\boldsymbol{e}_{2},...,\boldsymbol{e}_{K}].$ (2)
This enables the model to interact with all the text embeddings of the referenced concepts.
III-B2 Audio Embedding
We follow the baseline models [2] by applying a pre-trained VAE model to encode the reference audio clip from the mel-spectrogram into intermediate representations in a latent space. Similar to the text embedding for reference captions, the latent representation for each referenced audio $\boldsymbol{a}_{k}$ is obtained from the VAE encoder $\textit{f}_{\text{VAE}}(\cdot)$, as follows:
$\boldsymbol{r}_{k}=\textit{f}_{\text{VAE}}(\boldsymbol{a}_{k})$ (3)
The embeddings $\boldsymbol{r}_{k},k=1,...,K,$ are then concatenated as the third input:
$\boldsymbol{R}=[\boldsymbol{r}_{1},\boldsymbol{r}_{2},...,\boldsymbol{r}_{K}]$ (4)
The VAE decoder is trained to turn this latent vector $\boldsymbol{r}_{k}$ to its original form $\boldsymbol{a}_{k}$, which is applied in the final stage in Section III-D to reconstruct the output waveforms.
Figure 3: The details of the MRC UNet network, where the model takes the reference feature $\boldsymbol{R}$ and $\boldsymbol{E}$, prompt feature $\boldsymbol{C}$ and the current noise data $\boldsymbol{z}_{\lambda}$ as inputs to generate the dynamics for denoised data $\boldsymbol{z}_{1}$ on step $\lambda$. The output $\mu(\cdot)$ can then be used for both training and inference.
III-C Audio Feature Generation
III-C1 Module Strategy (Rectified Flow Matching)
Traditional diffusion-based generative models, such as DDPMs [11], generate samples through iterative denoising from a Gaussian prior, requiring hundreds of steps to gradually turn the noise distribution into the target data. However, such approaches are highly dependent on noise schedules and can be slow in sampling. Rectified Flow Matching (RFM) [74, 75], on the other hand, addresses the issue of slow sampling and increases stability by learning a continuous flow between noise and data distributions.
Unlike DDPMs that operate on discrete time steps $t\in\{0,1,...,N\}$, RFM models introduce a continuous flow variable $\lambda\in[0,1]$ that smoothly interpolates between Gaussian noise $\boldsymbol{z}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and the target data representation $\boldsymbol{z}_{1}$. By replacing discrete time steps with a continuous flow parameter $\lambda$, RFM enables smooth noise-to-data interpolation and a stable, schedule-free training objective based on a constant velocity field. The noisy data $\boldsymbol{z}_{\lambda}$ at the flow parameter of $\lambda$ is defined by a linear equation:
$\boldsymbol{z}_{\lambda}=(1-\beta\lambda)\boldsymbol{z}_{0}+\lambda\boldsymbol{z}_{1}$ (5)
where $\beta=(1-\sigma)$ is a small positive constant (e.g., $\sigma=1\times 10^{-5}$) introduced to avoid degeneracy and to improve numerical stability. This formulation allows the model to define a continuous transformation path between the noise and the target.
To train the model, a neural network $\mu(\cdot)$, which will be introduced as MRC in the next section, is optimized to predict the velocity field $\boldsymbol{v}=\boldsymbol{z}_{1}-\beta\boldsymbol{z}_{0}$. Unlike denoising targets in general diffusion models, this velocity field is independent of $\lambda$, which provides a consistent supervision signal in all flow positions sampled and improves training stability. The network is trained to minimize the following objective:
$L_{\text{RFM}}(\theta)=\mathbb{E}_{\lambda,\boldsymbol{z}_{1},\boldsymbol{z}_{0}}\left\|\mu(\boldsymbol{z}_{\lambda},\boldsymbol{R},\lambda,\boldsymbol{E},\boldsymbol{C})-\boldsymbol{v}\right\|^{2}$ (6)
Here, the network is conditioned on multiple modalities, including the reference audio features $\boldsymbol{R}$, the text embedding $\boldsymbol{E}$, prompt $\boldsymbol{C}$ and the continuous flow parameter $\lambda$, thus enabling fine-grained control over the generation.
During inference, an ODE solver integrates the dynamics defined by the trained network $\mu(\cdot)$, starting from Gaussian noise $\boldsymbol{z}_{0}$ and producing $\boldsymbol{z}_{1}$. Compared to the hundreds or even thousands of iterations required in diffusion-based sampling [10], this process typically requires fewer than 50 integration steps to synthesize high-quality audio.
III-C2 Module Structure (Multi-Reference Customization)
To generate the customized audio feature with referenced concepts in a tuning-free manner, a multi-reference customization structure is designed for the generator module. As illustrated in Figure 3, the module integrates both text and audio feature from input prompt and reference concepts and predicts the dynamics to generate the denoised output $\boldsymbol{z}_{1}$. Our module is inspired by ControlNet [30], where a similar structure is designed to connect the reference feature and the generated outputs. In MRC, a UNet structure with new encoder blocks during down-sampling to handle referenced concepts, and external connections between new encoder blocks and decoder blocks to integrate features during up-sampling. Different from ControlNet, two groups of encoder blocks are designed with same structure and two groups of the reference feature are connected separately with decoder blocks and cross-attention blocks respectively.
Specifically, the proposed down-sampling process consists of two groups of encoder paths: feature encoding path for the generated feature vector and referenced feature extraction path for the reference feature vector. In the down-sampling stages, the encoder layers of the feature encoding take the noisy latent feature $\boldsymbol{z}_{\lambda}$ and the text embedding $\boldsymbol{C}$. The encoder layers of referenced feature extraction, on the other hand, take the latent representation of the referenced audio concept $\boldsymbol{R}$ and the corresponding textual embedding $\boldsymbol{E}$. This approach allows DreamAudio to identify and extract the audio feature for the customized content based on the reference audio and their corresponding captions. It is noted that two groups of encoder blocks share the same structure, e.g., convolutional layers with a cross-attention module for text embedding, but are applied with independent weights for different downsampling purposes, respectively.
For the up-sampling stages, we feed all the features and conditions through a group of decoder blocks, designed as the feature decoding path. Specifically, the feature of the generated output and the extracted audio feature are integrated during the up-sampling stages. The reference audio features from each scaling stage are then considered as external conditions forwarded to the upsample layer within the same level. Different from single text embedding conditions in the feature encoding path, both the target prompt $\boldsymbol{E}$ and reference captions $\boldsymbol{R}$ are given as conditions through two cross-attention blocks, respectively.
III-D Audio Feature Reconstruction
As shown in the structure of Figure 2, DreamAudio leverages both the VAE decoder and a GAN-based vocoder for reconstructing the audio feature from latent space into the target waveform. Followed by the RFM model, the generated audio representation is first decoded into the mel-spectrogram by the VAE decoder. In the next stage, a vocoder is applied to convert the audio feature into the waveform as the final output, and we train the vocoder using the SOTA structure BigVGAN [50] in the proposed system.
III-E Customized Data Processing
In the early training stages, we observed that providing excessively detailed audio features caused the model to become overly reliant on these given representations, thereby limiting its ability to independently learn and generate new audio characteristics. To mitigate this, augmentation strategies are implemented to enhance the generalisation ability beyond provided references and improve robustness. Firstly, we randomly mask the contents of the customised audio event for the target audio. Secondly, we randomly remove the customised content during training. We train the DreamAudio model with a $10\%$ content masking and a $40\%$ content dropping rate. Experiments of the effectiveness of these masking and dropping metrics are introduced in Section VI-D3.
IV Datasets and Evaluation Benchmark
Previous text-to-audio generation models [2, 52, 15] mainly work on public audio-language datasets [12, 13, 14], which contains audio clips and their corresponding captions. In customized generation tasks, additional reference concepts, such as audio clips and their corresponding captions are also required. To this end, we construct the customized dataset for training and evaluation of the proposed model based on four commonly used audio-language datasets. The construction of these datasets enables DreamAudio to extract the customized feature from specific user contents. In addition, we manually collect a small-scale dataset with real-world customized scenarios as the first benchmark for CTTA. In the following sections, we first introduce the datasets collected as the base dataset to form the customized datasets. Then, we discuss the strategies for developing the training and testing data, followed by the introduction of the benchmark dataset. All evaluation benchmarks used for the system are presented in the end.
Figure 4: The generation pipeline of the customized datasets, with the Customized-Concatenation on the left and Customized-Overlay on the right. All the clips are selected randomly from the base dataset and both the concatenation clips and overlapped clips are fixed into 10-seconds.
TABLE I: The setup of four datasets used for training and testing of the proposed DreamAudio.The reference-type indicates the strategy for generating the reference concept and the reference-num presents the number of audio events for customization through the reference concept.
Dataset General-Dataset Customized-content Train-Num Test-Num Reference-Type Referenced-Num
AudioCaps-General AudioCaps $49,502$ $928$ Retrieval-based
Customized-Overlay WC+UB8K+ESC $146,481$ $200$ Overlap-audio 131-3
Customized-Concatenation AC+WC+UB8K+ESC $92,299$ $200$ Concate-audio 151-5
Customized-Fantasy Online-Collect - $25$ Concate-audio 33
IV-A General Datasets
IV-A1 AudioCaps
AudioCaps (AC) [31] is one of the largest audio datasets with hand-crafted captions. As a subset of the AudioSet [12], AudioCaps contains $52,905$ $10$-second audio clips, each sample is matched with a human-annotated caption. Taking from original videos on YouTube, AudioCaps contains a variety of classes of audio, including music, Foley sounds [55] and human speech. Each audio clip has a single caption in the training set and five captions in the test set, where we only apply the first caption of each waveform to the testing data. We follow the official dataset split and collect $49,502$ audio clips to construct the training datasets and $928$ clips for building the testing datasets.
IV-A2 WavsCaps
WavCaps (WC) [13] is a machine-labeled dataset of audio captions generated through Large Language Models (LLMs). WavCaps contains $403,050$ audio clips collected from various datasets such as AudioSet [12] and FreeSound [76], providing various durations of audio clips ranging from $0.1$s seconds to $68$ seconds. To align the length of the input and target audio clips, we only collect the samples with length of less than $10$ seconds, resulting in a group of $163,818$ audio samples. Then, we randomly select $95\%$ of the data for training and $5\%$ for testing.
IV-A3 UrbannSound8K
UrbanSound8K (UB8K) [77] contains $8,732$ labeled audio samples and each clip provides less than $4$ seconds of sound. The dataset is divided into $10$ different classes, including urban noise, background sound sources and natural sound sources. We randomly selected $8,000$ samples for training and $732$ for testing.
IV-A4 ESC-50
ESC-50 (ESC) [78] dataset has $2,000$ 55-second audio recordings which are evenly categorized into $50$ distinct classes including natural sounds and domestic sounds. For the split of training and testing data, we randomly choose $1,600$ samples in ESC-50 for training and $400$ samples for testing.
IV-B Customized Datasets
Based on the four general datasets mentioned in the previous section, we developed three different pipelines to generate customized datasets for training and evaluation. An overview of each dataset is presented in Table I and the pipelines for generating such datasets are illustrated in Figure 4.
IV-B1 Customized-Concatenation
We develop the dataset by concatenating audio events. In detail, we collect short audio clips of less than $5$ seconds from the database and then randomly form a group of audio clips which lengths sum up to $10$ seconds. The target audio is then generated by concatenating these short audio clips. Next, we construct the referenced audio by grouping the short audio clips into three sets, each concatenated into a single clip as the reference audio. For the reference captions, we first transfer the labels from US8K and ESC-50 into short captions by simply adding verbs and subjects to provide more understandable captions. Then, we generate the sentence by connecting the caption of each short audio clip with a connection phrase randomly selected from our connection list. For the Customized-Concatenation dataset, we collect $92,300$ audio-text pairs for training and $200$ pairs for testing.
IV-B2 Customized-Overlay
We find that the simple concatenating strategy led to several limitations. First, not all audio events can be connected smoothly, which may leave blank space. Second, audio created solely by concatenation often sounds unnatural, failing to accurately reproduce user-provided reference content, thus not effectively reflecting realistic scenarios. To guide the model with more realistic cases, we develop a second version of the customized dataset, called Customized-Overlay dataset, by mimicing real-world scenarios. Specifically, for each audio sample, we take a $10$-second audio clip as the base waveform and randomly select two shorter clips (less than $5$ seconds) as the “front” and “back” audio clips. The target audio clip is then generated by adding the “front” and “back” audio events into the base waveform. We remix the waveform of extended clips and base clips under a random signal to noise ratio (SNR) between -15 and 15 dB. The captioning strategy is similar to the approach we apply in the Customized-Concatenation dataset. We generated $146,481$ pairs for training and $200$ pairs for testing in the Customized-Overlay dataset.
IV-B3 Customized-Fantasy
We manually create a small-scale dataset as a benchmark for this new task, named as Customized-Fantasy. All the data clips are collected from online sources, including many unique and special events extracted from games or movies to simulate the user cases for real-world audio-production scenarios. For example, customization in laser guns, monster roars and Minion speaks. In addition, human interactions are involved to ensure that the captions are meaningful and present semantic coherence. In total, we collected 60 different events from Pixabay  [^1] https://pixabay.com/, a website which provides royalty-free audio resources under the Pixabay License suitable for research purposes, and compose $25$ distinct scenarios for zero-shot CTTA tasks.
IV-B4 AudioCaps-General
To maintain the capability of the system in more robust situations, e.g., prompting on audio events without the reference concepts. We apply the general AudioCaps dataset with empty reference concepts. For instance, the model is given with the target audio clip and the target caption, along with three empty audio clips and empty text as the reference concept. The same training-testing split is applied for training and evaluation of the model.
IV-C Evaluation Metrics
We use three different metrics for performance evaluation, including the Fréchet Audio Distance (FAD), Kullback-Leibler (KL) divergence and Contrastive Language-Pretraining (CLAP) score, by following the evaluation protocol of baseline audio generation models. In addition, we follow the evaluation pipelines from the customized image generations [22, 69] by incorporating subjective assessments that measure the consistency and the fidelity between generated and target audio, including CLAP audio score (CLAPA), AudioBox Aesthetics and subjective evaluations. Specifically, generative-based metrics including FAD, KL and CLAP-score are applied for evaluating the performance on Customized-Concatenation, Customized-Overlay and AudioCaps, subjective evaluations and AudioBox Aesthetics are used for the evaluations on Customized-Fantasy.
IV-C1 FAD
FAD first computes the multivariate Gaussian of two groups of embedding values from a pre-trained VGGish [79] feature extraction model. Then, it computes the Frechet distance between the Gaussian mean and variance of two sets of high-dimensional features. A lower FAD score indicates that the generated audio group presents a closer data distribution to the target audio group based on audio features.
IV-C2 KL Divergence
KL measures the logarithmic difference between the probabilities assigned by the distributions of two audio clips across all possible events. We apply the audio classification model PANNs [80] for the feature extraction and calculate the KL score between each target audio sample and their corresponding outputs. A lower KL presents a closer distance on distributions, suggesting the generated outputs are more similar to the target clips.
TABLE II: The setup of the primary experiments we performed, with DDPM for using DDPM-based training strategies and UNet for basic diffusion-based structure. DreamAudio with dual-encoder and single-encoder indicates whether the two groups of encoder blocks share the same weights.
Model Param(train) Diffusion MRC
DreamAudio-DDPM 760760M DDPM Dual-Encoder
DreamAudio-UNet 760760M RFM
DreamAudio-SEncoder 815815M RFM Single-Encoder
DreamAudio 279279M RFM Dual-Encoder
DreamAudio-L 891891M RFM Dual-Encoder
IV-C3 CLAP
The CLAP score calculates the cosine similarity between the text embedding and the audio embedding provided by the CLAP model [51]. The CLAP model learns projectors to align the audio and text embeddings into a joint space and presents paired audio and language data with similar features within the latent space. Given both audio embedding ${e}_{a}$ and text embedding ${e}_{t}$, the score is calculated as:
$\text{CLAPScore}(x,r)=\frac{{e}_{a}\cdot{e}_{t}}{\text{max}(\|{e}_{a}\|\|{e}_{t}\|,\epsilon)},$} (7)
where $\epsilon$ is a small value to avoid zero division. The CLAP score illustrates the correspondence between the generated audio and text prompt, and a higher score shows better performance on the semantic level.
IV-C4 CLAPA
Instead of calculating the similarity between audio feature and text feature, the CLAPA quantifies the relevance between the target audio feature embedding ${e}_{a}$ and the generated audio feature embedding $^{e}_{\hat{a}}$, by calculating their similarity as follows:
$\text{CLAPScore}_{A}(x,r)=\frac{{e}_{a}\cdot{e}_{\hat{a}}}{\text{max}(\|{e}_{a}\|\|{e}_{\hat{a}}\|,\epsilon)}.$} (8)
A higher score means that the generated output is more similar to the target audio within the feature space.
IV-C5 Subjective Evaluation
We adopt the evaluation metrics used in the baseline models [21], incorporating both Overall Impression (OVL) and Audio-Text Relation (REL) for subjective assessments. Detailed scoring instructions, including illustrative examples, are provided to the raters to ensure clarity. For OVL, we follow the approach outlined in the baseline models [21] by asking raters: How would you rate the overall quality of this audio sample?. Responses are scored on a Likert scale, ranging from 5 for “excellent” to 1 for “bad”. For REL, we adapt the metric to better suit the customized generation tasks by asking: Does the generated audio successfully present the target audio events customized by the reference audio samples? The subjective evaluations were performed on ten different listeners, with six researchers in audio and speech and four from other fields.
IV-C6 AudioBox Aesthetics
AudioBox Aesthetics is a recently proposed audio quality assessment tool that has demonstrated a strong correlation with human-based subjective evaluations [81]. In detail, this metric applies a transformer-based model for extracting and assessing the quality of audio clips without reference clips. In this paper, we choose the production quality (PQ) and content usefulness (CU) for the task. In detail, PQ is a technical objective aspect of the overall quality and CU is a subjective axis to evaluate the likelihood of audio sample for content creation.
TABLE III: Comparison of model performances on the customized-based evaluation set. AC is short for AudioCap dataset, CM is for Customized-Concatenation and CE is for Customized-Overlay.
Model Dataset Customized-Concatenation Customized-Overlay Customized-Fantasy
[…]
Re-AudioLDM AudioCaps $3.05$ $3.24$ $39.4$ $48.7$ $2.96$ $3.09$ $44.3$ $47.8$ $5.80$ $5.05$ $3.15$ $3.26$
DreamAudio-DDPM AC+CM+CE $0.94$ $1.09$ $52.3$ $79.8$ $0.99$ $1.32$ $\mathbf{46.7}$ $83.3$ $6.18$ $5.45$ $3.70$ $3.53$
DreamAudio-UNet AC+CM+CE $1.59$ $1.73$ $48.1$ $73.5$ $1.03$ $0.87$ $41.8$ $77.8$ $5.59$ $4.78$ $3.41$ $3.39$
DreamAudio-SEncoder AC+CM+CE $0.79$ $1.05$ $51.4$ $85.9$ $0.88$ $0.92$ $41.5$ $81.3$ $6.12$ $5.31$ $3.51$ $3.69$
DreamAudio AC+CM+CE $0.50$ $0.90$ $52.0$ $86.3$ $0.78$ $0.82$ $42.0$ $81.7$ $6.31$ $5.56$ $3.74$ $3.91$
DreamAudio-L AC+CM+CE $\mathbf{0.46}$ $\mathbf{0.92}$ $\mathbf{52.5}$ $\mathbf{87.7}$ $\mathbf{0.73}$ $\mathbf{0.67}$ $42.4$ $\mathbf{83.9}$ $\mathbf{6.37}$ $\mathbf{5.65}$ $\mathbf{3.89}$ $\mathbf{4.17}$
V Experiments Details
V-A Model Architecture Details
For the text encoder, we use a pre-trained Flan-T5 model [53], resulting in a $1024$ dimension feature sequence with a fixed length of $50$ for every caption. We use the pre-trained VAE model from AudioLDM which provides a compression ratio of $4$ and results in a latent vector of $256$ dimension in temporal and $16$ dimension in frequency for a ten-second mel spectrogram. For the vocoder, we apply BigVGAN [50] and perform the self-supervised pre-training on $16$ kHz audio clips. Table II summarizes our experiments. We perform the experiments with two sizes of the latent diffusion model, DreamAudio and DreamAudio-L, with the hidden dimension size of $n_{\text{hidden}}=96$ and $n_{\text{hidden}}=128$, respectively. In addition, we performed evaluations of several variants of the proposed system. The DreamAudio system trained with DDPM strategies is denoted as DreamAudio-DDPM. The DreamAudio system with general UNet structure is named as DreamAudio-UNet. The system with shared-weight MRC structure is denoted as DreamAudio-SEncoder, where the encoder blocks for referenced feature extraction and feature encoding share the same weights.
V-B Training and Inference Setup
Despite the pre-trained text encoder and VAE, we train the model and vocoder separately. We first train the vocoder on a large-scale audio dataset for 11M steps and then freeze the module into the proposed DreamAudio to train the model. Combining the whole system, we train the model with two customized datasets mentioned in Section IV-B. The system is trained on a single NVIDIA A100 80GB GPU for $2$ millions steps. Following similar strategies from generative networks, we apply the Classifier Free Guidance (CFG) for RFM with the value of $2.0$ for both DreamAudio and DreamAudio-L. We utilize the AdamW [82] optimizer with a learning rate of $5\times 10^{-5}$ for training and apply a linear warming up for $10,000$ steps.
VI Results and Analysis
We evaluated our proposed system on both customized and general text-to-audio generation. The following section first discusses the performance of our system on the CTTA tasks and then explores the capability of general TTA tasks.
VI-A Customized Text-to-Audio Generation
All the models in Table III are evaluated on both Customized-Concatenation and Customized-Overlay datasets. We compare our result with Re-AudioLDM [20] as this is the only work related to customized and zero-shot tasks. The result for AudioBox-TTA-RAG [29] is not included because the authors did not provide any model or checkpoint. For Re-AudioLDM, we apply the system with the retrieval number of $3$ as the retrieval-based information. As shown in Table III, the proposed DreamAudio significantly outperforms the previous systems across most of the metrics. The baseline system, Re-AudioLDM, achieves a FAD score of $3.05$ and $2.96$ for Customized-Concatenation and Customized-Overlay, respectively. Our systems, on the other hand, illustrate a substantially enhanced score of $0.46$ for the concatenated testing set and $0.73$ for the overlapped testing set. DreamAudio-L also attains the best KL divergence score of $0.92$ and $0.67$, considerably improved the performance upon any previous methods.
For the CLAP score, our system demonstrates SOTA performance on the Customized-Concatenation dataset and achieves comparable results on the Customized-Overlay dataset. Unlike other objective metrics that directly compare target and generated audio samples, the CLAP score measures semantic alignment rather than precisely evaluating the presence of customized audio content. In this case, it is less effective in evaluating the task that focuses on specific customization details. In contrast, the CLAPA score evaluates the similarity between the features from each generated audio clip and its corresponding target audio clip, making it a more accurate metric for reflecting the correctness of customized content. DreamAudio-L achieves a CLAPA score of $87.7$ on the Customized-Concatenation dataset and $83.9$ on the Customized-Overlay dataset, significantly surpassing all previous models. For the Customized-Fantasy testing set, DreamAudio achieves the best performance among production quality, content usefulness and human evaluation, where DreamAudio-L achieves SOTA results on OVL of $3.89$ and SOTA performance on REL of $4.17$. The significant enhancement offered by DreamAudio in both OVL and REL illustrates that the proposed system can generate specific audio features related to the reference concepts, as well as being faithful to text prompts.
TABLE IV: Comparison of model performances on the AudioCaps evaluation set. AC is short for AudioCaps dataset, CM is for Customized-Concatenation and CE is for Customized-Overlay. The dataset marked with $\ast$ indicates fine-tuning on AudioCaps.
[…]
AudioLDM AC+AS+2 others $5.25$ $1.90$ $42.1$ $53.5$
AudioGen AC+AS+8 others $2.87$ $1.52$ $46.4$ $60.2$
Make-an-Audio AC+AS+13 others $2.39$ $1.64$ $45.4$ $59.8$
Tango AudioCaps $2.24$ $1.04$ $51.7$ $66.2$
AudioLDM2 AC+AS+6 others $2.56$ $1.75$ $45.8$ $55.5$
Re-AudioLDM AudioCaps $1.85$ $1.46$ $49.9$ $62.0$
DreamAudio AC+CM+CE $4.25$ $2.48$ $34.9$ $43.6$
DreamAudio AudioCaps $1.92$ $1.51$ $47.5$ $58.8$
VI-B General Text-to-Audio Generation
In this section, we evaluate the capability of DreamAudio on general text-to-audio generation tasks. We compare the performance on audio generation with several SOTA systems, including AudioLDM [2], AudioGen [16], Make-an-Audio [19], Tango2 [52], AudioLDM2 [21], and Re-AudioLDM [20]. All the checkpoints are downloaded from open-sourced GitHub [^2] https://github.com/ or HuggingFace [^3] https://huggingface.co/ to reproduce the results. As shown in Table IV, these results demonstrate that our DreamAudio can maintain good performance in general text-to-audio tasks even without access to the corresponding customized content. For instance, DreamAudio fine-tuned on the AudioCaps dataset (providing empty reference contents) achieves a FAD score of $1.92$ and a KL score of $1.51$, which outperforms all non-retrieval methods. Compared to current SOTA models, DreamAudio achieves competitive CLAP scores of $47.5$ and $58.8$ in terms of semantic and audio feature similarity, respectively. DreamAudio did not outperform the baseline models probably because the MRC module did not provide useful audio information for general tasks. As a result, the model loses the advantage of incorporating external features and instead relies primarily on its own capability for independent audio feature generation.
Figure 5: The details of the MRC UNet network for reference length fine-tuning. All the previous blocks are frozen and only the new introduced CNN layer is trained during this stage.
TABLE V: The experiment on customized generation with four reference concepts, models are fine-tuned based the model with reference number of $3$, step number with indicates the model is directly applied without any further fine-tuning .
[…]
3 $0^{\ast}$ $0.50$ $92$ $52.5$ $87.7$
4 $0^{\ast}$ $14.31$ $6.01$ $34.9$ $43.6$
4 $10,000$ $3.86$ $3.01$ $45.6$ $69.5$
4 $50,000$ $0.79$ $1.15$ $49.2$ $77.5$
VI-C Customized Generation on Various Reference Number
In this section, experiments are carried out to verify the capability of DreamAudio on various referenced concept pairs. To fit for different lengths of the referenced VAE latent, an extra CNN alignment layer is applied to squeeze the referenced feature so the concatenated feature being forwarded to the decoder blocks is designed as a fixed size, e.g. $96$ in the temporal dimension and $6$ in the frequency dimension. As depicted in Figure 5, we freeze all the blocks except the alignment blocks, which are trained when DreamAudio is fine-tuned for different lengths of reference concepts. The results of the proposed model for four reference pairs are presented in Table V. The experiment illustrates that DreamAudio can adapt the extended reference information within only $10,000$ steps and achieve a competitive FAD score of $0.79$ in only $50,000$ steps, which is comparable to the $0.50$ FAD score of DreamAudio on three pairs of reference concepts.
VI-D Ablation Studies
In order to validate our design choice of the proposed system, a series of ablation experiments were conducted on each proposed technique within DreamAudio.
VI-D1 Effectiveness of RFM
As shown in Table III, we compare the outputs of DreamAudio with different training approaches on various datasets. Experiments on RFM-based systems present SOTA performance on customized tasks, i.e., Customized-Concatenation and Customized-Overlay evaluations. On the other hand, we observe that DreamAudio-DDPM trained with diffusion-based techniques offers better results in the semantic feature than DreamAudio, giving better CLAP scores. These experimental results demonstrate two key findings: a) RFM-based models show promising performance in generating customized and personalized content; b) DDPM-based approaches guide the models toward more robust performance by preserving the capability of generating features based on semantic information.
VI-D2 Effectiveness of MRC
We conduct various experiments with and without the UNet blocks with MRC in Table III For DreamAudio-UNet, the input dimension is extended four times to undertake both denoised latent and reference latent to align the process of MRC. The results illustrate that the proposed system with basic UNet structures exhibits significant performance degradation among all the metrics. In addition, the comparison between DreamAudio-SEncoder and DreamAudio indicates that the use of two groups of encoder blocks for encoding and feature extraction helps improve performance.
TABLE VI: Ablation Studies on the data masking, data dropping and the effectiveness of different proportions of training data. The general task is evaluated on the AudioCaps testing set and CTTA is evaluated on the Customized-Concatenation and Customized-Overlay testing sets. For data processing, AC is short for AudioCaps, and $C$ and CO are short for Customized-Concatenation and Customized-Overlay. The number indicates the quantity of each dataset for training.
Model Training Data-Mixing General Text-to-Audio Task Customized Text-to-Audio Task
[…]
DreamAudio 4949K 9292K 146146K $10\%$ $40\%$ $4.25$ $2.48$ $34.9$ $43.6$ $0.64$ $0.86$ $47.0$ $84.0$
$\% 00\%$ $5.28$ $2.65$ $31.2$ $40.4$ $\mathbf{0.51}$ $\mathbf{0.56}$ $47.7$ $\mathbf{88.1}$
$\% 1010\%$ $4.28$ $2.25$ $35.2$ $44.4$ $0.56$ $0.71$ $47.2$ $84.3$
$50\%$ $90\%$ $3.81$ $2.15$ $38.1$ $46.9$ $1.32$ $1.68$ $42.5$ $74.3$
2424K 9292K 146146K $10\%$ $40\%$ $5.58$ $2.69$ $30.1$ $39.2$ $0.64$ $0.88$ $47.2$ $84.9$
4949K 2424K 2424K 2.85 1.82 41.5 51.1 $2.11$ $1.96$ $41.9$ $68.5$
4949K 9292K 2424K $5.14$ $2.68$ $33.9$ $42.1$ $0.48$ $0.61$ $\mathbf{47.8}$ $86.5$
4949K 2424K 146146K $4.03$ $2.11$ $36.8$ $44.6$ $0.88$ $1.01$ $44.5$ $80.2$
VI-D3 Effectiveness of Data Masking and Dropping
We also conducted experiments on reference concept masking, dropping, and the impact of different data proportions. As shown in Table VI, the masking and dropping ratios significantly influence the model’s performance on customized tasks. When provided with more precise reference concepts, the model learns more effectively. However, based on the KL divergence and CLAPA scores, we observe that the generated results contain more similar information to the target audio. This may be because the model is led to pay attention to extraction and concatenation rather than generation, which may explain the reason of performance drop in general TTA generation tasks. This indicates that without masking and dropping, the model’s performance is degraded. Furthermore, in the data distribution comparison, we found that an excessive amount of customized-concatenation data makes the model overly reliant on the content of the reference concept. In contrast, Customized-Overlay data helps maintain a balance between leveraging the reference concept and the text prompts. Training with AudioCaps data improves the robustness of the model in general TTA scenarios. However, an excessive amount of AudioCaps data causes the model to shift focus away from the features provided by the reference concept, leading to degraded performance on the CTTA task.
VI-E Limitation
Despite giving the SOTA performance, our method has certain limitations.
VI-E1 Fixed Reference Format
The requirement for reference concept on both customized audio and related captions can be challenging to satisfy in real-world applications. Preparing suitable captions for referenced audio samples can lead to extra workloads and pose a practical barrier to general users.
VI-E2 The Number of References
The current model architecture limits the number of reference concept pairs to three, restricting the system’s flexibility when handling extensive or diverse customization. Ablation studies show that DreamAudio can be easily fine-tuned to handle a varying number of reference concepts. However, the performance of the system is dependent on the availability of appropriate training data.
VI-E3 Artificial Training Data
Due to the lack of real-world data for customized tasks, the audio generated by the proposed model can sometimes sound unnatural. Addressing this limitation requires the development of more diverse and high-quality datasets tailored to specific tasks, allowing models to capture nuanced patterns and improve their generalization to practical applications.
VII Conclusion and Future Works
In this paper, we have introduced DreamAudio, an audio generation model for customized text-to-audio generation (CTTA). Given target prompts and reference audio-caption pairs as input, DreamAudio demonstrates SOTA performance in generating audio clips with specified audio contents in the CTTA task, while also delivering competitive results in the general TTA task, establishing a robust foundation for a wide range of audio-related applications. To further validate the real-world applicability of our method, we created the first benchmark for CTTA, which includes zero-shot audio events to simulate real-world scenarios. Our evaluation demonstrates that DreamAudio captures these user-specific audio events, highlighting its ability to generate meaningful and semantically coherent content. In addition, we studied the adaptability of the proposed method, which can effectively handle a varying number of audio-caption pairs as reference concepts. Future work will focus on improving the flexibility of incorporating reference audio-text pairs, changing the number of references, and developing diverse and high-quality real-world data. We will also extend the model to the integration of additional modalities such as images or video, and its use in other customized audio tasks, such as audio separation, style transfer, and editing.
Acknowledgments
This research was partly supported by a research scholarship from the China Scholarship Council (CSC), an internship from ByteDance, funded by British Broadcasting Corporation Research and Development (BBC R&D), Engineering and Physical Sciences Research Council (EPSRC) Grant EP/T019751/1 “AI for Sound”, and a PhD scholarship from the Centre for Vision, Speech and Signal Processing (CVSSP), University of Surrey. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising.